LBC Technical Observations 8 April 2010 UT

07:30 Starting LBC observing. LUCIFER technical observing was done during the first half

07:32 Focus/Collimate on ACT0294 at 83 deg elevation, FPIA crashed - coma in blue. We had to enter total Z7=-2500nm and Z8=+1000nm to SX.

07:49 Blue pupil images suddenly have no stars. Is shutter stuck closed? We login to lbccontrol to power cycle the shutter. This worked.

08:02 ACT0294 observed for IE and CA... 80135
Red Pointing Center is 1080, 2914. Changed IE and CA: CA=-21.7 and IE=-38. (IA is currently -267). Temperature is -0.26 (glass) and we are at 80 deg elevation.

08:11 ACT090 observed for IA. 24 deg elevation... no significant change in IA this time. T=-0.4 ambient and mirror.

08:18 Fast OB to slew to Mellote 111 --- this is an open cluster with many stars for FPIA work. elevation 66 degrees. Collimate/focus with RedLimit=BlueLimit=200nm. Take extra/intra focal pairs: R 083532 083621 B 083535 083624

08:47 BlueZ22Zero was 2.3. We now changed this in lbcfpia.cfg to BlueZ22Zero = 1.5. Now focussing/collimating with again limits=200. Temperatures of glass are similar to ambient.

09:00 extrafocal has loads of negative Z11. We are going to change BlueFinalZ11Zero (which was -300). Image pairs are: B 090103/090152 and R 090100/090149

09:15 Taking a series of extra-foc images while adjusting Z11. 091506 - first Z11 = +500nm 091753 still needs positive Z11 - Z11 = +500nm again 092048 this looks flat

09:23 Changed BlueFinalZ11Zero to +700 (was -300). After "convergence" we ran extra/intra focal pair. B 093338/093427 R 093336/093425 We see a little bit of Z22, so we changed BlueZ22Zero from 1.5 to 1.6. Exit and restart IDL

09:38 Collimating/focussing again. extra/intra B 094534/094625. There is Z11 in these.

09:51 Previous field at 12 h is getting low (50 deg elev) and so we are slewing to Dzim 5 (16:27, +38). Elevation 79 deg.
Run dofpia to focus/collimate (limits=200 as before). Take extra/intra focal images: 100522/100611 Extra has large inner hole - Send -300nm Z11 to SX and take extra/intra: 101246/101336 still the extra has a large hole Send -300nm Z11 to SX again (+700 - 600). Ambient T=0.1, SXglass=-0.3, DXglass=-0.4 101615/101707 now extra has slight ring around central hole and intra has slight ring around outer... Send +500 while running dofpia...
So - we think we will need to add some Z11 (so -400 instead of -600).

10:24 Finished dofpia, Send -400 Z11 to SX extra/intra focal pupils: 102558/102637 still extra has a slight ring around central hole... still Z11 is too negative. Add +100 nm Z11 to SX (-300 after dofpia). 103006/103055 relatively flat, though hole sizes still are too big/too small for extra(60pxl outer)/intra(52.8pxl outer) Adjusted BlueFinalZ11Zero to +400 (from +700).

Restarted IDL with this new BlueFinalZ11Zero = +400. To summarize, we changed BlueFinalZ11Zero=+400 and BlueZ22Zero=1.6

10:35 Fine tuning matrix
*Z22=+300nm:*
Run dofpia and then send +300 Z22 manually. Run dofpia again to analyze results (we changed Z4,Z11,Z22 gains to 0 so no corrections will be sent for these... no, we attempted this but apparently TCS_SendWavefront sets gains to 1).
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  2627    336     27    182    -11  -1000   -100    y  0.83  1.22                      /newdata/lbcb.20100408.104334.fits  V-BESSEL
 -2364   -131   -110   -318     -4    600    -57    y  0.90  1.02                      /newdata/lbcb.20100408.104506.fits  V-BESSEL
A factor -9 on focus... but cannot tell about Z11.

*Z22 = +100nm*
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1785    203    144    188   -127   -680    -51    y  1.02  0.84                      /newdata/lbcb.20100408.105755.fits  V-BESSEL
 -1033    -12    332   -188    -33    115      5    y  0.85  0.44                      /newdata/lbcb.20100408.105924.fits  V-BESSEL
   674    142   -249     12     65   -188    -21    y  0.81  0.31                      /newdata/lbcb.20100408.110053.fits  V-BESSEL
 -229    -17    159     -2    -83      0    -20    y  0.78  0.13                      /newdata/lbcb.20100408.110217.fits  V-BESSEL
  -228   -177    -41      8     66    151    -21    y  1.27  0.16                      /newdata/lbcb.20100408.110339.fits  V-BESSEL
    82      0      0      0      0    -43      0    y  0.00  0.00                     dofpia_sph3_adjustment             

11:13 Running dofpia and when converged, Andrew will give Z11 = +500nm to both SX and DX --- looks good.

11:17 going to measure filter focus offsets. Run rV superfoc sequence ..112040.. V-BESSEL best focus is 0.0 r-SLOAN best focus 0.027

RUs superfoc sequence Us best focus = 0.40 but not a good fit ("W") R-BESSEL best focus = 0.017

11:30 We were running dofpia again prior to taking another rV superfoc but we found that the Z4 values were oscillating. Still the ambient and glass temps are close (0.1 ambient and glass = 0.1 SX and -0.22 DX). Stopping and restarting dofpia.

One more focus sequence at rV before it gets too light... B 120148, R 120145
V-BESSEL is not a convincing fit but is this time at -0.03. r-SLOAN best focus at 0.015mm

B-BESSEL best focus is 0.044 z:

One more at rV ... B ... 121724... , R ...121707... V-BESSEL -0.029

~12:25 Too bright for any more dofpia

Slew to flat field position to take a flat...Us Y
12:30 Run OB with no preset (dec = -90). Does it dither?? B 123217... and R 123214... no --- RA and DEC are fixed. Pause/stop and run OB with preset. This OB does cause the telescope to dither. So as expected the preset is now needed. Will have to make a script to input coords to generic flat field OBs.

12:45 10 biases with dome closed & lights off.

end of night

-- OlgaKuhn - 08 Apr 2010
Topic revision: r5 - 10 Apr 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback